# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 3.1. Problem Setup
In this section, we consider the same setup as [19], i.e., we focus on the PDE with the following form.
@@
where $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable. In the following, we assume that $\|\bm{x}\|_{2}\leq 1$ for $\bm{x}\in[0,T]\times\bar{\Omega}$ and $f,g$ are bounded continuous functions.
@@
where $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}\in\mathbb{R}^{m(d+2)}$, $\bm{a}=(a_{1},\cdots,a_{m})^{T}\in\mathbb{R}^{m}$ and for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+2}$ is the weight vector of the first layer, $a_{r}$ is the output weight and $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function. Similar as that in Section 2, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we aim to minimize the following empirical loss function.
$L(\bm{w},\bm{a})$ $:=\sum\limits_{p=1}^{n_{1}}\frac{1}{2n_{1}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w},\bm{a})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w},\bm{a})-f(\bm{x}_{p})\right)^{2}$ (15)
$\quad+\sum\limits_{j=1}^{n_{2}}\frac{1}{2n_{2}}\left(\phi(\bm{y}_{j};\bm{w},\bm{a})-g(\bm{y}_{j})\right)^{2}.$
Similar as that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights. Then the gradient descent updates the hidden weights by the following formulations:
$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k),\bm{a})}{\partial\bm{w}_{r}}$ (16)
for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. For brevity, we write $L(\bm{w})$ for $L(\bm{w},\bm{a})$.
To simplify the notations, for the residuals of interior and boundary, we denote them by $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ respectively, i.e.,
$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial u}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (17)
@@
$\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ (20)
and
$\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}.$ (21)
At this time, we have
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\sum\limits_{p=1}^{n_{1}}s_{p}(\bm{w})\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(\bm{w})\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$ (22)
and the Gram matrix $\bm{H}(\bm{w})$ is defined as
$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\ \bm{D}:=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right].$ (23)
3.2. Main Results
First, we make the following assumptions about the training samples, which are similar to those for the $L^{2}$ regression problems.
Assumption 3.
Assumption 4.
Under Assumption 4, Lemma 3.3 in [19] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$. Similarly, $\bm{H}^{\infty}$ plays an important role in the optimization process. The following two lemmas indicate that $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$, leading to the linear convergence.
Lemma 4.
Lemma 5.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-(n_{1}+n_{2})^{2}e^{-mR}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
$\|H(\bm{w})-H(0)\|_{F}<CM^{2}R,$ (24)
where $M=2(d+2)\log(2m(d+2)/\delta)$ and $C$ is a universal constant.
In [19], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\left(\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right),$ (25)
which leads to the requirement $\eta=\mathcal{O}(\lambda_{0})$. Instead, we can generalize easily the method used in the $L^{2}$ regression problems to PINNs and obtain the following recursion formula.
Lemma 6.
For all $k\in\mathbb{N}$, we have
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k),$ (26)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
for $j\in[n_{2}]$,
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
By using the recursion formula, we arrive at our main result.
Theorem 2.
Under Assumption 3 and Assumption 4, if we set the number of hidden nodes
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{bmatrix}\bm{s}(0)\\ \bm{h}(0)\end{bmatrix}\right\|_{2}^{2}$ (29)
for all $k\in\mathbb{N}$.
Remark 3.
It may be confused that [19] has uesd the same method in [10], but it only requires $\eta=\mathcal{O}(\lambda_{0})$. Actually, it is due to that the loss function of PINN has been normalized. If we let $n_{1}=n_{2}=n$ and $\widetilde{\bm{H}}^{\infty}$ be the Gram matrix induced by unnormalized loss function of PINN, then $\lambda_{min}(\bm{H}^{\infty})=\lambda_{min}(\widetilde{\bm{H}}^{\infty})/n$, leading to the convergence rate similar to that of regression problem. Due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ is independent of the sample size $n$, but may depends on the dimension $d$.
4. Convergence of Natural Gradient Descent for Two-Layer Physics-Informed Neural Networks
4.1. Problem Setup
Although we have improved the learning rates of gradient descent for $L^{2}$ regression problems and PINNs, one needs to set the learning rates to be small enough due to the agnostic $\|\bm{H}^{\infty}\|_{2}$. For instance, we may need to set it to be $\mathcal{O}(1/n)$ for the $L^{2}$ regression problem. Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which may be slow with small $\lambda_{0}$. [21] and [16] have provided the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer neural networks for $L^{2}$ regression problems. They showed that the maximal learning rate can be $\mathcal{O}(1)$ and the convergence rate is independent of $\lambda_{0}$, which result in a faster convergence rate. In the section, we conduct the convergence analysis of NGD for PINNs and show that NGD indeed leads to an even faster convergence rate.
We consider the same setup as Section 3 and aim to minimize the following empirical loss function via NGD.
$L(\bm{w}):=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right).$ (30)
The NGD gives the following update rule:
$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$ (31)
where $\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$ is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate. Specifically, for $p\in[n_{1}]$,
@@ 4.1. Problem Setup
we consider both settings that $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function or a smooth activation function that satisfies the following assumption. From Lemma 3.3 in [19] and Lemma 2 in [23], we know that $\bm{H}^{\infty}$ is strictly positive definite in both settings and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
Assumption 5.
There exists a constant $c>0$ such that $|\sigma(0)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
@@
where $k\in\{0,1,2,3\}$. Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function.
Unlike the approach for gradient descent, [21] and [16] focus on the change of the Jacobian matrix for NGD, rather than focusing on the Gram matrix. More precisely, they show that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$. However, this approach does not work for PINNs. Instead, we consider the stability of $\bm{J}(\bm{w})$ with respect to every single weight vector $\bm{w}_{r}$.
Lemma 7 (Stable Jacobian).
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-P(\delta,m,R)$ the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satsify for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
@@
$P(\delta,m,R)=1-\delta-n_{1}e^{-mR};$ (37)
(2) when $\sigma(\cdot)$ satisfies Assumption 5, we have that
@@
With the stability of Jacobian matrix, we can derive the following main result.
Theorem 3 (Convergence of NGD).
@@
(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 4, we set
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
@@
(2) When $\sigma(\cdot)$ satisfies Assumption 5, under Assumption 4, we set
$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$

### v3

@@ 2. Problem Setup
In this section, we consider the same setup as [21], focusing on the PDE with the following form.
@@
where $\Omega\subset\mathbb{R}^{d}$ is an open and bounded domain, $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable. In the following, we assume that $\|\bm{x}\|_{2}\leq 1$ for $\bm{x}\in[0,T]\times\bar{\Omega}$ and $f,g$ are bounded continuous functions.
@@
where $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}\in\mathbb{R}^{m(d+2)}$, $\bm{a}=(a_{1},\cdots,a_{m})^{T}\in\mathbb{R}^{m}$ and for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+2}$ is the weight vector of the first layer, $a_{r}$ is the output weight and $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function. Here, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
Similar to that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights. In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we denote $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ by
$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (3)
@@
where $\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ and $\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}$.
The gradient descent updates the hidden weights by the following formulations:
$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}_{r}}$ (6)
for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. The Gram matrix $\bm{H}(\bm{w})$ is defined as $\bm{H}(\bm{w})=\bm{D}^{T}\bm{D}$, where
$\bm{D}:=\left(\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right).$ (7)
3. Improved Results of GD for Two-Layer PINNs
Assumption 3.1.
Assumption 3.2.
Under Assumption 3.2, Lemma 3.3 in [21] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\bm{H}(\bm{w})]$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$. Similar to the case of the regression problem in [10], $\bm{H}^{\infty}$ plays an important role in the optimization process. Specifically, under over-parameterization and random initialization, we have two facts that (1) at initialization $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and (2) for any iteration $k\in\mathbb{N}$, $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$. The following two lemmas can be used to verify these two facts, which are crucial in the convergence analysis.
Lemma 3.3.
Remark 3.4.
Under the premise of deriving the same conclusion as Lemma 3.3, Lemma 3.5 in [21] requires that $m=\tilde{\Omega}\left(\frac{(n_{1}+n_{2})^{4}}{(n_{1}n_{2})^{2}\lambda_{0}^{2}}\left(\log(\frac{1}{\delta})\right)^{7}\right)$, where some terms involving $\log(m)$ are omitted. In contrast, on one hand, our conclusion is independent of $n_{1}$ and $n_{2}$, and on the other hand, our conclusion exhibits a clear dependence on $d$. Moreover, the method in [21] involves truncating the Gaussian distribution and then applying Hoeffding’s inequality, which is quite complicated. In contrast, we utilize the concentration inequality for sub-Weibull random variables, which serves as a simple framework for this class of problems.
Lemma 3.5.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-n_{1}e^{-mR}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<CM^{2}R,$ (8)
where $M=2(d+2)\log(2m(d+2)/\delta)$ and $C$ is a universal constant.
Remark 3.6.
Lemma 3.6 in [21] shows that when $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}\leq R=\tilde{\mathcal{O}}\left(\frac{\lambda_{0}\delta}{(n_{1}+n_{2})(\log m)^{3}}\right)$ holds for all $r\ in[m]$, then $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4}$. In contrast, Lemma 3.5 only requires $R=\mathcal{O}\left(\frac{\lambda_{0}}{d^{2}(\log(m/\delta)^{2}}\right)$ to reach same result.
For the $L^{2}$ regression problem, as shown in [10], the convergence of gradient descent requires that the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, where $n$ is the sample size of the regression problem. It is evident that this requirement on the learning rate is difficult to satisfy in practical scenarios, since $\lambda_{0}$ is unknown and $n^{2}$ is too large . For PINNs, [21] follows the methodology of [10], thus inheriting similarly stringent requirements on the learning rate. By investigating a new decomposition method for the residual, we arrive at our main result.
Theorem 3.7.
Under Assumption 3.1 and Assumption 3.2, if we set the number of hidden nodes
$m=\Omega\left(\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}$ (9)
for all $k\in\mathbb{N}$.
Remark 3.8.
It may be confusing that [21] has used the same method in [10], yet it only requires $\eta=\mathcal{O}(\lambda_{0})$. Actually, it is because that the loss function of PINN has been normalized. If we let $n_{1}=n_{2}=n$ and $\widetilde{\bm{H}}^{\infty}$ be the Gram matrix induced by unnormalized loss function of PINN, then $\lambda_{min}(\bm{H}^{\infty})=\lambda_{min}(\widetilde{\bm{H}}^{\infty})/n$, leading to the convergence rate similar to that of regression problem. At this point, due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ can be bounded by the trace of $\bm{H}^{\infty}$, which is an explicit constant that is independent of the sample size $n_{1},n_{2}$. Therefore, in practice, we can set the learning rate to satisfy the theoretical convergence requirement, bridging the gap between theory and practice.
Remark 3.9.
Regarding the impact of dimensionality on the convergence of GD for PINNs, Theorem 3.7 consists of two parts: one is explicit, namely $d^{8}$, and the other is implicit, specifically $\lambda_{0}^{4}$, whose relationship with dimensionality remains unclear. The explicit impact arises from the form of the PDE. For instance, the PDE (1) we consider contains $\mathcal{O}(d)$ terms. In concurrent work, [26] investigated the impact of dimensionality and the order of PDEs on convergence under the setting of continuous gradient flow. Specifically, for PDEs of order 2, the form they considered includes $\mathcal{O}(d^{2})$ terms, and the explicit dependence on dimensionality is $d^{28}$. On the one hand, [26] only addressed the continuous gradient flow case, while the discrete case requires more refined analysis as stated in [10]. On the other hand, our results can naturally extend to the PDEs they considered, where the explicit dependence on dimensionality is $d^{16}$, which is better than the result in [26]. Investigating the lower bounds of the smallest eigenvalue of the NTK for PINNs, similar to what has been done for deep ReLU neural networks in [27], represents a promising direction for future research.
4. Convergence of NGD for Two-Layer PINNs
Although we have improved the learning rate of gradient descent for PINNs, it may still be necessary to set the learning rates to be sufficiently small for some complex PDEs. Because, although for all PDEs, $\bm{tr}(\bm{H}^{\infty})$ is an explicit constant, it depends on the form of the PDE, and for complex PDEs, it may be quite large. Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which depends on the sample size and may be extremely small. [19] and [20] have provided the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer neural networks for $L^{2}$ regression problems. They showed that the maximal learning rate can be $\mathcal{O}(1)$ and the convergence rate is independent of $\lambda_{0}$, which result in a faster convergence rate. However, their methods cannot generalize directly to PINNs. In the section, we conduct the convergence analysis of NGD for PINNs and demonstrate that it results in a faster convergence rate for PINNs compared to gradient descent.
In this section, we consider the same setup as described in Section 2. Specifically, we focus on the PDE of the form given in (1) and follow the same initialization as described in Section 2. During the training process, we fix the output weight $\bm{a}$ and update the hidden weights via NGD. The optimization objective is the empirical loss function presented in (5), which is defined as follows:
The NGD gives the following update rule:
$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},$ (11)
@@ 4. Convergence of NGD for Two-Layer PINNs
$\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$
is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate. Specifically, for $p\in[n_{1}]$,
@@
Remark 4.1.
We note that [19] and [20] have independently and concurrently established the convergence of NGD in the context of regression problems. The difference lies in the fact that [19] focused on ReLU activation functions, whereas [20] considered smooth activation functions and consistently set the learning rate to $1$. Here, following [19], we refer to this approach as NGD. In [20], the authors derived this method based on NTK kernel regression and termed it the Gram-Gauss-Newton (GGN) method.
Remark 4.2.
The classical Gauss-Newton method is given by
$\bm{w}(k+1)=\bm{w}(k)-\left(\bm{J}(k)^{T}\bm{J}(k)\right)^{-1}\bm{J}(k)^{T}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k),\end{pmatrix}$
which is different from our NGD method (11). The Gram matrix $\bm{J}(k)^{T}\bm{J}(k)$ of classical Gauss-Newton method will be extremely high-dimensional and singular for large network parameters, while our NGD $\bm{J}(k)\bm{J}(k)^{T}$ won’t.
@@
we consider settings where $\sigma(\cdot)$ is either the $\text{ReLU}^{3}$ activation function or a smooth activation function satisfying the following assumption.
Assumption 4.3.
There exists a constant $c>0$ such that $\sup_{z\in\mathbb{R}}|\sigma^{(3)}(z)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
@@
where $k\in\{0,1,2,3\}$. Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function. We also assume that for any positive integer $n\geq 2$, $\lim\limits_{x\to+\infty}\sigma^{(n)}(x)/\phi(x)=c_{n}\neq 0$, where the function $\phi(\cdot)$ satisfies that
$\lim\limits_{x\to+\infty}\phi(x)=0,\lim\limits_{x\to+\infty}x\cdot\frac{\phi(bx)}{\phi(x)}=0$ (16)
holds for any constant $b>1$.
Lemma 4.4.
If no two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}\cup\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel, then the Gram matrix $\bm{H}^{\infty}$ is strictly positive definite for activation functions that satisfy Assumption 4.2, i.e., $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$.
Remark 4.5.
Assumption 4.3 holds for various commonly used activation functions, including logistic function $\sigma(z)=1/(1+e^{-z})$ (with $\phi(z)=e^{-z}$), softplus function $\sigma(z)=\log(1+e^{z})$ (with $\phi(z)=e^{-z}$), hyperbolic tangent function $\sigma(z)=(e^{z}-e^{-z})/(e^{z}+e^{-z})$ (with $\phi(z)=e^{-2z}$), swish function $\sigma(z)=z/(1+e^{-z})$ (with $\phi(z)=ze^{-z}$) and others. Compared to the $\text{ReLU}^{3}$ activation function, these smooth activation functions are more popular for PINNs because solving PDEs typically requires high-order derivatives.
Unlike the approach for gradient descent, [19] focus on the change of the Jacobian matrix for NGD rather than the Gram matrix. More precisely, they demonstrate that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$. Roughly speaking, they show that when $\|\bm{w}-\bm{w}(0)\|_{2}$ is small, then $\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}$ is also proportionately small. However, this approach is not applicable to PINNs, because the loss function involves derivatives. Roughly speaking, the stability considered in [19] is more global in nature, whereas ours is local. Since the subsequent conclusions require the boundedness of local weights, we do not use this stability. Moreover, from Theorem 1 in [19], we can see that this stability imposes additional constraints on the learning rate. Therefore, we instead focus on the stability of $\bm{J}(\bm{w})$ with respect to each individual weight vector $\bm{w}_{r}$, which provides a more targeted approach.
Lemma 4.6.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-P(\delta,m,R)$ the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
@@
$P(\delta,m,R)=\delta+n_{1}e^{-mR};$ (18)
(2) when $\sigma(\cdot)$ satisfies Assumption 4.3, we have that
@@
With the stability of Jacobian matrix, we can derive the following convergence results.
Theorem 4.7.
@@
(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 3.2, we set
$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
@@
(2) When $\sigma(\cdot)$ satisfies Assumption 4.3, under Assumption 3.2, we set
$m=\Omega\left(\frac{1}{1-\eta}\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
@@
In Theorem 4.7, the requirements of $m$ with $\text{ReLU}^{3}$ and smooth activation functions exhibit different dependencies on $\lambda_{0}$ and $d$. The discrepancy is primarily due to the distinct formulations presented in (16) and (18) of Lemma 4.5.
Remark 4.8.
We first compare our results with those of NGD for $L^{2}$ regression problems. Given that the convergence results are the same, our focus shifts to examining the necessary conditions for the width $m$. As demonstrated in [19] and [20], it is required that $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}\delta^{3}}\right)$ for ReLU activation function and $m=\Omega\left(\max\left\{\frac{n^{4}}{\lambda_{0}^{4}},\frac{n^{2}d\log(n/\delta)}{\lambda_{0}^{2}}\right\}\right)$ for smooth activation function. Clearly, our result has a worse dependence on $d$, which is inevitable due to the involvement of derivatives in the loss function. Moreover, our requirement for $m$ appears to be almost independent of $n$, primarily because our loss function has been normalized. With smooth activation functions, in addition to the dependence on $d$, Theorem 4.7 (2) only requires that $m=\Omega(\lambda_{0}^{-3})$. However, [20] demands a more stringent condition, requiring that $m=\Omega(\lambda_{0}^{-4})$.
Continuing our analysis, we contrast our results with those of GD for PINNs. Roughly speaking, [21] has shown that when $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, $m=\widetilde{\Omega}\left(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}}\right)$ and $\eta=\mathcal{O}(\lambda_{0})$, then the convergence result (9) holds. It is evident that our result, i.e, Theorem 4.7 (1), has a milder dependence on $n_{1},n_{2}$ and $\delta$. Furthermore, the learning rate and convergence rate are independent of $\lambda_{0}$, resulting in faster convergence.
Comparing with our results in Section 3, the requirement for $m$ in Theorem 4.7 (1) is the same as in Theorem 3.8, when we make $\eta$ less close to $1$. On the other hand, since $\eta=\mathcal{O}(1)$ and the convergence rate only depends on $\eta$, NGD can lead to faster convergence than GD.